# Unified Stochastic Framework for Neural Network Quantization and Pruning

### v2

@@ 1. Introduction
In this theory-focused paper, we extend post-training algorithms for quantization and pruning by building on the work presented in [29], which introduced Stochastic Path Following Quantization (SPFQ). SPFQ is a stochastic error-correcting algorithm that sequentially quantizes a neuron’s weights to minimize accumulated error, with an accompanying theoretical upper bound on the error that scales as $\sqrt{\log(N)}$, where $N$ is the dimension of the neuron. This scaling compares favorably to the $\sqrt{N}$ scaling associated with other techniques. However, SPFQ’s focus on minimizing accumulated error without constraints limits its applicability to extremely low-bit quantizers, as it cannot guarantee rigorous error bounds in such cases. Additionally, while SPFQ includes an analysis for finite-bit quantization when the number of bits is not too small, its framework does not readily extend to pruning.
@@ 3. Stochastic Path Following Compression
Parameters: Stochastic operator $\mathcal{T}$, failure threshold $\theta>0$, constant $C\geq 1$.
@@
$q=\textsc{Compression}(\mathcal{T},\theta,C,w,X,\widetilde{X}).$
@@
Algorithm 1 uses $m$ data points in $\mathbb{R}^{N_{0}}$ represented by $X\in\mathbb{R}^{m\times N_{0}}$ to guide the compression. The basic idea is to perform the compression one layer at a time, starting with the first. Thus, at layer $i$, we have access to $X^{(i-1)}:=\Phi^{(i-1)}(X)$ and $\widetilde{X}^{(i-1)}:=\widetilde{\Phi}^{(i-1)}(X)$, both in $\mathbb{R}^{m\times N_{i-1}}$, and the goal is to compress $W^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ into $Q^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ such that $X^{(i-1)}W^{(i)}\approx\widetilde{X}^{(i-1)}Q^{(i)}$. To accomplish this we use Algorithm 2 to independently, and possibly in parallel, replace each column (neuron) $w\in\mathbb{R}^{N_{i-1}}$ of $W^{(i)}$, entry by entry, with a new neuron $q$ such that $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$. In Algorithm 2, to ensure $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$, we proceed sequentially by computing, for $t=1,2,...,N_{i-1}$, the quantized or pruned coefficient $q_{t}$ that aligns the sums $\sum_{j=1}^{t}w_{j}X^{(i-1)}_{j}$ and $\sum_{j=1}^{t}q_{j}\widetilde{X}^{(i-1)}_{j}$. This step forms the core of our method. It balances two competing requirements, namely reducing the error while ensuring that $q_{t}$ lies within the quantization alphabet or has a high likelihood of being set to zero, depending on the application.
@@
In the case of quantization, if $\tfrac{\langle h_{t},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}$ falls outside of the boundaries of the chosen alphabet, the error can become uncontrollable. In [29], where $C=1$, this is prevented by simply adding more elements to the alphabet. In contrast, in order to work with 1-bit alphabets we allow $C\geq 1$, set a threshold $\theta$, and terminate our algorithm if $\left|\tfrac{\langle u_{t-1},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}\right|>\theta$. For instance, if $\{-2K,2K\}$ is the alphabet, then $\theta$ is set to $K$. Meanwhile, for pruning, no such termination is necessary, so $\theta$ is set to infinity. In the remainder of this paper, we show that these modifications allow us to control the error with high probability on the stochasticity of the operator $\mathcal{T}$.
Algorithm 2 Compression($\mathcal{T},\theta,C,w,X,\widetilde{X}$)
Parameters: Stochastic operator $\mathcal{T}$, failure threshold $\theta>0$, constant $C\geq 1$.
@@
3:     if $\left|\tfrac{\langle u_{t-1},\widetilde{X}_{t}\rangle}{C\|\widetilde{X}_{t}\|^{2}}\right|>\theta$ then
4:         Algorithm Fails
5:         return
6:     end if
7:     Compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$
8:     Compute $v_{t}=\tfrac{\langle h_{t},\widetilde{X}_{t}\rangle}{C\|\widetilde{X}_{t}\|^{2}}$
9:     Compute $q_{t}=\mathcal{T}(v_{t})$
10:     Update accumulated error: $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}\widetilde{X}_{t}$
11:end for
@@ 4.1. Main Theorem
We now define convex ordering, a notion that will allow us to control random vectors generated by our algorithm by replacing them with Gaussian random vectors.
@@
In order to state and prove our main result on the error associated with Algorithm 2 when $X=\widetilde{X}$, we will first introduce an equivalent formulation of our algorithm. Recall that we initialize the accumulated error as $u_{0}=0$ and note that at each step $t$, we first check a failure condition. If the condition is not met, we compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$. Then we use $h_{t}$ to compute $v_{t}=\tfrac{\langle h_{t},X_{t}\rangle}{C\|X_{t}\|^{2}}$ and compute the compressed entry $q_{t}=\mathcal{T}(v_{t})$. Finally, we update the accumulated error $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}X_{t}$ and move on to the next step. Since the algorithm fails when the failure condition is met at any step, we also have to track whether the algorithm fails at some step. Thus, instead of studying $u_{t}$ directly, we introduce and analyze an the auxiliary quantity $\widetilde{u}_{t}$, which encodes the information about failure. The sequence $\widetilde{u}_{t}$ with $\widetilde{u}_{1}=u_{1}$ is recursively defined via
$\widetilde{u}_{t}$ =$=\begin{cases}0,&\text{if }\left|\frac{\langle\widetilde{u}_{i-1},X_{i}\rangle}{C\|X_{i}\|^{2}}\right|>\theta\text{ for any }i\leq t,\\ \widetilde{u}_{t-1}+w_{t}X_{t}-q_{t}X_{t},&\text{otherwise}.\end{cases}$
[…]
Using convex ordering, we can control the probability $\mathbb{P}\{\text{$|\tfrac{\langle\widetilde{u}_{i-1},X_{i}\rangle}{C\|X_{i}\|^{2}}|\leq\theta$ for all $i$ and }\|\widetilde{u}_{N}\|_{\infty}\leq\kappa\}$ by showing that $\widetilde{u}_{i}$ is dominated by a Gaussian random vector. This result, stated in the theorem below, allows us to control the errors and probabilities for the quantization and pruning applications in Sections 5, 6, and 7. Here we adapt the proof technique from [1].
@@
Let $C\geq 1$. Assume the weight vector satisfies $\|w\|_{\infty}\leq K$. Further, assume that the operator $\mathcal{T}$ satisfies $\mathbb{E}[\mathcal{T}(v)]=v$ and $|\mathcal{T}(v)-v|\leq M$ for all $v$ with $|v|\leq K+\theta$. Then, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, the resulting $\widetilde{u}_{j}$ satisfies
$\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\beta_{j}I),\quad\text{for }1\leq j\leq N,$
where $\beta_{j}$ is given by
$\beta_{j}=\frac{C\pi M^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2}.$
@@
First by Lemma 4.4, we deduce $\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\Sigma_{j})$, where $\Sigma_{j}$ is defined as in Lemma 4.4. Further by Lemma 4.5, we have $\Sigma_{j}\preceq\beta_{j}I$. We then apply Lemmas 4.3, c and a to complete the proof. ∎
@@ 4.2. Technical Lemmas
    If $X\prec_{cx}Y$ and $Y\prec_{cx}Z$, then $X\prec_{cx}Z$.
@@
    If $X\prec_{cx}Y$, then for any linear transformation $M$ on $\mathbb{R}^{n}$, we have $MX\prec_{cx}MY$.
@@
    If $A$ and $B$ are two positive semi-definite matrices and $A\preceq B$, then $\mathcal{N}(0,A)\prec_{cx}\mathcal{N}(0,B)$.
@@
    Let $X$, $Y$, $W$, and $Z$ be random vectors and suppose that $X$ and $Y$ live on the same probability space. Let $W$ and $Z$ be independent and suppose that $X\prec_{\mathrm{cx}}W$ and $(Y-X)|X\prec_{\mathrm{cx}}Z$. Then $Y\prec_{\mathrm{cx}}W+Z$.
@@
    Let $X$ be a real-valued random variable with $\mathbb{E}X=0$ and $|X|\leq C$. Then $X\prec_{\mathrm{cx}}\mathcal{N}\bigl{(}0,\frac{\pi C^{2}}{2}\bigr{)}$.
@@
    Let $X$ be a symmetric random vector in $\mathbb{R}^{n}$, i.e., $X$ and $-X$ have the same distribution, and let $A\subset\mathbb{R}^{n}$ be a symmetric set (i.e., $x\in A\Rightarrow-x\in A$). Define $X^{\prime}$ by
    $X^{\prime}=\begin{cases}X,&\text{if }X\in A^{c},\\ 0,&\text{otherwise.}\end{cases}$
    Then $X^{\prime}\prec_{cx}X$.
  7. g.
@@
Using Lemma 4.3 allows to deduce the following result on the distribution of $\widetilde{u}_{j}$. We use a proof technique similar to that in [1].
@@
Under the same assumptions as in Theorem 4.2, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, we obtain $\widetilde{u}_{j}$ such that:
$\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\Sigma_{j})\text{ for $1\leq j\leq N$},$
where $\Sigma_{0}=0$ and $\Sigma_{j},j\geq 1$ is defined recursively as
$\Sigma_{j}=\Big{(}I-\dfrac{P_{X_{j}}}{C}\Big{)}\Sigma_{j-1}\Big{(}I-\dfrac{P_{X_{j}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{j}X_{j}^{\top}.$
@@
We use induction. For $j=0$, the statement is trivial. Assume the statement is true for $j=t-1$. Define $u^{\prime}_{t-1}$ as
$u^{\prime}_{t-1}=\left\{\begin{split}&\widetilde{u}_{t-1},&\text{ if $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$,}\\ &0,&\text{ otherwise.}\end{split}\right.\quad$
The random variable $\widetilde{u}_{t-1}$ is symmetric, so $u^{\prime}_{t-1}\prec_{cx}\widetilde{u}_{t-1}$ by Lemma 4.3, f. If $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$, then the algorithm does not fail and we have
$\widetilde{u}_{t}$ $=\widetilde{u}_{t-1}+w_{t}X_{t}-q_{t}X_{t}$
$=\widetilde{u}_{t-1}+w_{t}X_{t}-\widetilde{v}_{t}X_{t}+(\widetilde{v}_{t}-q_{t})X_{t}$
$=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\widetilde{u}_{t-1}+(\widetilde{v}_{t}-\mathcal{T}(\widetilde{v}_{t}))X_{t}$
due to the definition of $\widetilde{v}_{t}$ and $q_{t}$. If $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|>\theta$, then we have $\widetilde{u}_{t}=0$ by definition. Thus we can express $\widetilde{u}_{t}$ using $u^{\prime}_{t-1}$ as
$\widetilde{u}_{t}=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u^{\prime}_{t-1}+R_{t}X_{t}$
where
$R_{t}=\left\{\begin{split}&\widetilde{v}_{t}-\mathcal{T}(\widetilde{v}_{t}),&\text{ if $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$,}\\ &0,&\text{ else.}\end{split}\right.\quad$
When $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$, we have
$|\widetilde{v}_{t}|=\Big{|}\dfrac{\langle\widetilde{h}_{t},X_{t}\rangle}{C\|X_{t}\|^{2}}\Big{|}=\Big{|}w_{t}+\dfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}\Big{|}\leq K+\theta.$
By the assumption about $\mathcal{T}$, $R_{t}$ is mean zero and $|R_{t}|\leq M$. We then use e from Lemma 4.3, to deduce $R_{t}\prec_{cx}\mathcal{N}(0,\tfrac{\pi M^{2}}{2})$. Further applying Lemma 4.3, b, we obtain
$R_{t}X_{t}\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
Since $u^{\prime}_{t-1}\prec_{cx}\widetilde{u}_{t-1}\prec_{cx}\mathcal{N}(0,\Sigma_{t-1})$, we have
$\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u^{\prime}_{t-1}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Big{)}.$
Finally, by Lemma 4.3, d, we conclude
$\widetilde{u}_{t}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
@@
Having controlled our object of interest by a Gaussian random vector with a known covariance matrix, $\Sigma_{j}$ defined in Lemma 4.4, we next control the covariance matrix itself.
@@
Let $C\geq 1$, then for the above defined $\Sigma_{j}$, we have
$\Sigma_{j}\preceq\beta_{j}I,$
where $\beta_{0}=0$ and $\beta_{j}=\dfrac{C\pi M^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2},$ when $j\geq 1$.
@@
The base case $j=0$ is obvious. Let us now proceed by induction. Suppose $\Sigma_{t-1}\preceq\beta_{t-1}I$. Then

### v3

@@ 1. Introduction
In this theory-focused paper, we extend post-training algorithms for quantization and pruning by building on the work presented in [29], which introduced Stochastic Path Following Quantization (SPFQ). SPFQ is a stochastic error-correcting algorithm that sequentially quantizes a neuron’s weights to minimize accumulated error, with an accompanying theoretical upper bound on the Euclidean error that scales as $\sqrt{\min\{m,N\}}\cdot\text{polylog}(N)$, where $N$ is the dimension of the neuron and $m$ is the number of calibration data points. This scaling compares favorably to the $N$ scaling associated with the most common Round-to-Nearest (RTN) method. However, SPFQ’s focus on minimizing accumulated error without constraints limits its applicability to extremely low-bit quantizers, as it cannot guarantee rigorous error bounds in such cases. Additionally, while SPFQ includes an analysis for finite-bit quantization when the number of bits is not too small, its framework does not readily extend to pruning.
@@ 3. Stochastic Path Following Compression
Parameters: Stochastic operator $\mathcal{T}$, constant $C\geq 1$.
@@
$q=\textsc{Compression}(\mathcal{T},C,w,X,\widetilde{X}).$
@@
Algorithm 1 uses $m$ data points in $\mathbb{R}^{N_{0}}$ represented by $X\in\mathbb{R}^{m\times N_{0}}$ to guide the compression. The basic idea is to perform the compression one layer at a time, starting with the first. Thus, at layer $i$, we have access to $X^{(i-1)}:=\Phi^{(i-1)}(X)$ and $\widetilde{X}^{(i-1)}:=\widetilde{\Phi}^{(i-1)}(X)$, both in $\mathbb{R}^{m\times N_{i-1}}$, and the goal is to compress $W^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ into $Q^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ such that $X^{(i-1)}W^{(i)}\approx\widetilde{X}^{(i-1)}Q^{(i)}$. To accomplish this, we use Algorithm 2 to independently, and possibly in parallel, replace each column (neuron) $w\in\mathbb{R}^{N_{i-1}}$ of $W^{(i)}$, entry by entry, with a new neuron $q$ such that $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$. In Algorithm 2, to ensure $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$, we proceed sequentially by computing, for $t=1,2,...,N_{i-1}$, the quantized or pruned coefficient $q_{t}$ that aligns the sums $\sum_{j=1}^{t}w_{j}X^{(i-1)}_{j}$ and $\sum_{j=1}^{t}q_{j}\widetilde{X}^{(i-1)}_{j}$. This step forms the core of our method. It balances two competing requirements, namely reducing the error while ensuring that $q_{t}$ lies within the quantization alphabet or has a high likelihood of being set to zero, depending on the application.
@@
In the case of quantization, if the ratio $\tfrac{\langle h_{t},\widetilde{X}^{(i-1)}{t}\rangle}{C|\widetilde{X}^{(i-1)}{t}|^{2}}$ falls outside the range of the chosen alphabet, the resulting error can become unbounded. In [29], where $C=1$, this issue is avoided by enlarging the alphabet to include more quantization levels. In contrast, to work with a fixed 1-bit alphabet, we allow a scaling constant $C\geq 1$ to ensure the quantization remains controlled. In Theorem 1, we show that choosing $C\sim\log(N_{0}N_{1})$, where $N_{0}$ and $N_{1}$ are the input and output channel dimensions respectively, guarantees that the quantized weights remain within the 1-bit alphabet, while still achieving a logarithmic reconstruction error bound on $\|XW-XQ\|_{\infty}$ with high probability..
In the remainder of this paper, we show that these modifications allow us to control the error with high probability on the stochasticity of the operator $\mathcal{T}$.
Algorithm 2 Compression($\mathcal{T},C,w,X,\widetilde{X}$)
Parameters: Stochastic operator $\mathcal{T}$, constant $C\geq 1$.
@@
3:     Compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$
4:     Compute $v_{t}=\tfrac{\langle h_{t},\widetilde{X}_{t}\rangle}{C\|\widetilde{X}_{t}\|^{2}}$
5:     Compute $q_{t}=\mathcal{T}(v_{t})$
6:     Update accumulated error: $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}\widetilde{X}_{t}$
7:end for
@@ 4.1. Main Theorem
We now define convex ordering [22], a notion that will allow us to control random vectors generated by our algorithm by replacing them with Gaussian random vectors.
@@
Recall that we initialize the accumulated error as $u_{0}=0$ and note that at each step $t$, we compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$. Then we use $h_{t}$ to compute $v_{t}=\tfrac{\langle h_{t},X_{t}\rangle}{C\|X_{t}\|^{2}}$ and compute the compressed entry $q_{t}=\mathcal{T}(v_{t})$. Finally, we update the accumulated error $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}X_{t}$ and move on to the next step.
Using convex ordering, we can prove that $u_{i}$ is dominated by a Gaussian random vector. This result, stated in the theorem below, allows us to control the errors and probabilities for the quantization and pruning applications in Sections 5, 6, and 7. Here we adapt the proof technique from [1].
@@
Let $C\geq 1$. Assume the weight vector satisfies $\|w\|_{\infty}<K$. Further, assume that the operator $\mathcal{T}$ satisfies $\mathbb{E}[\mathcal{T}(v)]=v$ and $|\mathcal{T}(v)-v|\leq M$ for all $v\in\mathbb{R}$. Then, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, the resulting $u_{t}$ satisfies
$u_{t}\prec_{cx}\mathcal{N}(0,\beta_{t}I),\quad\text{for }1\leq t\leq N,$
where $\beta_{t}$ is given by
$\beta_{t}=\frac{C\pi M^{2}}{2}\max_{1\leq i\leq t}\|X_{i}\|^{2}.$
@@
First by Lemma 4.4, we deduce $u_{t}\prec_{cx}\mathcal{N}(0,\Sigma_{t})$, where $\Sigma_{t}$ is defined as in Lemma 4.4. Further by Lemma 4.5, we have $\Sigma_{t}\preceq\beta_{t}I$. We then apply Lemmas 4.3, c and a to complete the proof. ∎
@@ 4.2. Technical Lemmas
    Let $X$, $Y$ and $Z$ be $n$-dimensional random vectors. If $X\prec_{cx}Y$ and $Y\prec_{cx}Z$, then $X\prec_{cx}Z$. (Lemma 2.3 in [1])
@@
    Let $X$, $Y$ be $n$-dimensional random vectors. If $X\prec_{cx}Y$, then for any matrix $M\in\mathbb{R}^{n\times n}$, we have $MX\prec_{cx}MY$. (Lemma A.3 in [29])
@@
    If $A$ and $B$ are two $n\times n$ positive semi-definite matrices and $A\preceq B$, then $\mathcal{N}(0,A)\prec_{cx}\mathcal{N}(0,B)$. (Lemma A.2 in [29])
@@
    Let $X$, $Y$, $W$, and $Z$ be $n$-dimensional random vectors and suppose that $X$ and $Y$ live on the same probability space. Let $W$ and $Z$ be independent and suppose that $X\prec_{\mathrm{cx}}W$ and $(Y-X)|X\prec_{\mathrm{cx}}Z$. Then $Y\prec_{\mathrm{cx}}W+Z$. (Lemma 2.5 in [1])
@@
    Let $X$ be a real-valued random variable with $\mathbb{E}X=0$ and $|X|\leq C$. Then $X\prec_{\mathrm{cx}}\mathcal{N}\bigl{(}0,\frac{\pi C^{2}}{2}\bigr{)}$. (Lemma 2.6 in [1])
@@
    (Lemma B.2 in [29])
Using Lemma 4.3 allows to deduce the following result on the distribution of $u_{j}$. We use a proof technique similar to that in [1].
@@
Under the same assumptions as in Theorem 4.2, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, we obtain $u_{t}$ satisfying
$u_{t}\prec_{cx}\mathcal{N}(0,\Sigma_{t})\text{ for $1\leq t\leq N$},$
where $\Sigma_{0}=0$ and $\Sigma_{t},t\geq 1$ is defined recursively as
$\Sigma_{t}=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}.$
@@
We use induction. For the base case, we have $u_{0}=0$, so the statement is trivially true. Now, assume $u_{t-1}\prec_{cx}\mathcal{N}(0,\Sigma_{t-1})$. By the update formula of $q_{t}$ in Algorithm 2:
$u_{t}$ $=u_{t-1}+w_{t}X_{t}-q_{t}X_{t}$
$=u_{t-1}+w_{t}X_{t}-v_{t}X_{t}+(v_{t}-q_{t})X_{t}$
$=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u_{t-1}+(v_{t}-\mathcal{T}(v_{t}))X_{t}.$
The third equality is due to the definition of $v_{t}$ and $q_{t}$. By the assumption on $\mathcal{T}$, and conditioning on $u_{t-1}$, $(v_{t}-\mathcal{T}(v_{t})\mid u_{t-1})$ is mean zero and $|v_{t}-\mathcal{T}(v_{t})|\leq M$. We then use e from Lemma 4.3, to deduce $(v_{t}-\mathcal{T}(v_{t})\mid u_{t-1})\prec_{cx}\mathcal{N}(0,\tfrac{\pi M^{2}}{2})$. Further applying Lemma 4.3, b, we obtain
$\left((v_{t}-\mathcal{T}(v_{t}))X_{t}\mid u_{t-1}\right)\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
Since $u_{t-1}\prec_{cx}\mathcal{N}(0,\Sigma_{t-1})$ by the induction hypothesis, we similarly have
$\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u_{t-1}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Big{)}.$
Finally, by Lemma 4.3, d with $\left(I-\frac{P_{X_{t}}}{C}\right)u_{t-1}$ in place of $X$, $\left(v_{t}-\mathcal{T}(v_{t})\right)X_{t}\mid u_{t-1}$ in place of $Y-X\mid X$ and $\mathcal{N}\left(0,\left(I-\frac{P_{X_{t}}}{C}\right)\Sigma_{t-1}\left(I-\tfrac{P_{X_{t}}}{C}\right)\right)$ in place of $W$, with $\mathcal{N}\left(0,\tfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\right)$ as $Z$, where $W$ and $Z$ are chosen independently, we conclude
$u_{t}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
@@
Having controlled our object of interest by a Gaussian random vector with a known covariance matrix, $\Sigma_{t}$ defined in Lemma 4.4, we next control the covariance matrix itself.
@@
Let $C\geq 1$, then for the above defined $\Sigma_{t}$, we have
$\Sigma_{t}\preceq\beta_{t}I,$
where $\beta_{0}=0$ and $\beta_{t}=\dfrac{C\pi M^{2}}{2}\max_{1\leq i\leq t}\|X_{i}\|^{2},$ when $t\geq 1$.
@@
The base case is obvious. Let us now proceed by induction. Suppose $\Sigma_{t-1}\preceq\beta_{t-1}I$. Then